OpenStudy (anonymous):

In baseball, the equation E=9R/I gives a pitchers run average E, where R is the number of earned runs the player allowed and I is the number of innings pitched. Last season, a pitcher had earned run average of 2.80 and allowed 70 earned runs. How many innings did the pitcher pitch last season?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[2.80 = \frac{ 9(70) }{ I }\]

OpenStudy (anonymous):

Haha I don't even know if I wrote this right o-o

OpenStudy (texaschic101):

looks good to me...can you finish it ?

OpenStudy (anonymous):

\[2.80 = \frac{ 630 }{ I }\]

OpenStudy (anonymous):

Then multiply by I?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

I(2.80) = 630

OpenStudy (texaschic101):

correct....now divide both sides by 2.80 to get I by itself

OpenStudy (anonymous):

Divide 630 by 2.80 o-o Fun xD

OpenStudy (texaschic101):

yep

OpenStudy (texaschic101):

so what did you get ?

OpenStudy (anonymous):

225

OpenStudy (anonymous):

I = 225?

OpenStudy (texaschic101):

you can even check your answer.. E = (9R) / I 2.80 = (9)(70) / 225 2.80 = 630/225 2.80 = 2.80 (correct) so I = 225

OpenStudy (anonymous):

Wow thanks @texaschic101 ;3

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Could you help me with another algebraic problem real quick?

OpenStudy (anonymous):

I just gotta find that paper o-o

OpenStudy (anonymous):

11) \[\frac{ 3 }{ 4 }t + \frac{ 5 }{ 6 } = \frac{ 2 }{ 3 }t \]